OpenStudy (anonymous):

Find the probability for the experiment of drawing two marbles (without replacement) from a bag containing three green, five yellow, and four red marbles such that both marbles are yellow. a. 5/33 b. 25/144 c. 25/132 d. 2/5 e. 5/7

OpenStudy (anonymous):

probability = good choices / total choices

OpenStudy (anonymous):

3G 5Y 4R probability of drawing yellow first time = 5/12 probability of drawing yellow again = 4/11 total probability = 5/12 . 4/ 11 = 5/33

OpenStudy (whpalmer4):

First marble has P=5/12 of being yellow. 2nd marble has P=4/11 of being yellow if first was yellow w/o replacement. Multiply the probabilities...